OpenStudy (rjtrain1):

Each month Terrance spends $128.00 on his car payment, $63.00 for car insurance, and $45.00 on gas. What is the exact amount of money Terrance spends each month to own a vehicle?

OpenStudy (francesca_marie):

$236

OpenStudy (rjtrain1):

thank you I also need help with Which addition property is being illustrated below? (Type commutative or associative) (6 + 7) + 1 = 6 + (7 + 1)

OpenStudy (francesca_marie):

okay hold on

OpenStudy (francesca_marie):

http://www.aaamath.com/pro74ax2.htm I think its Associative Property

OpenStudy (rjtrain1):

I aso need help with The what of 14 and 7 = 7

OpenStudy (francesca_marie):

The percentage?

OpenStudy (rjtrain1):

The options are product quotient difference and sum

OpenStudy (francesca_marie):

Sum

OpenStudy (rjtrain1):

thank you I also need help on Use the distributive property to evaluate the following expression: 9 × (4 + 9)

OpenStudy (francesca_marie):

Thats weird you are suppose to learn distributive property in 6th grade but I'll help anyways hold on

OpenStudy (rjtrain1):

I am in 6th grade not 5th miss said erlier

OpenStudy (francesca_marie):

Oh okay that are the options for the question?

OpenStudy (rjtrain1):

there is no options

OpenStudy (francesca_marie):

Oh okay its 117

OpenStudy (rjtrain1):

thank you one more for this question thin I will close it Which property is illustrated below? 1 + 7 = 7 + 1 identity property of addition commutative property of addition associative property of addition distributive property

OpenStudy (francesca_marie):

That's Commutative property if you need anymore help just tag me :)

OpenStudy (rjtrain1):

ok thank you I passed the quiz

OpenStudy (francesca_marie):

Thats great :)